Caspian Sea Drinks' is financed with 69.00% equity and the remainder in debt. They have 11.00-year, semi-annual pay, 5.76% coupon bonds which sell for 98.91% of par. Their stock currently has a market value of $24.53 and Mr. Bensen believes the market estimates that dividends will grow at 3.58% forever. Next year’s dividend is projected to be $2.33. Assuming a marginal tax rate of 34.00%, what is their WACC (weighted average cost of capital)?

Caspian Sea Drinks' is financed with 69.00% equity and the remainder in debt. They have 12.00-year, semi-annual pay, 5.30% coupon bonds which sell for 97.62% of par. Their stock currently has a market value of $24.14 and Mr. Bensen believes the market estimates that dividends will grow at 3.03% forever. Next year’s dividend is projected to be $3.00. Assuming a marginal tax rate of 23.00%, what is their WACC (weighted average cost of capital)?

Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))
Answer #1

1). kE = [D1 / P0] + g = [$2.33 / $24.53] + 0.0358 = 0.0950 + 0.0358 = 0.1308, or 13.08%

To find the kD, we need to put the following values in the financial calculator:

N = 11*2 = 22;

PV = -98.91%*1000 = -989.10

PMT = (5.76%/2) * 1000 = 28.80

FV = 1000;

Press CPT then I/Y, which gives us 2.95

So, Periodic Rate = 2.95%

kD = Periodic Rate * No. of compounding periods in a year = 2.95% * 2 = 5.90%

WACC = [wD * kD * 1 - t)] + [wE * kE]

= [0.31 * 5.90% * (1 - 0.34)] + [0.69 * 13.08%]

= 1.21% + 9.02% = 10.23%

2). kE = [D1 / P0] + g = [$3 / $24.14] + 0.0303 = 0.1243 + 0.0303 = 0.1546, or 15.46%

To find the kD, we need to put the following values in the financial calculator:

N = 12*2 = 24;

PV = -97.62%*1000 = -976.20

PMT = (5.30%/2) * 1000 = 26.50

FV = 1000;

Press CPT then I/Y, which gives us 2.79

So, Periodic Rate = 2.79%

kD = Periodic Rate * No. of compounding periods in a year = 2.79% * 2 = 5.57%

WACC = [wD * kD * 1 - t)] + [wE * kE]

= [0.31 * 5.57% * (1 - 0.23)] + [0.69 * 15.46%]

= 1.33% + 10.67% = 12.00%